OpenStudy (anonymous):

how do you find the derivatiove of a logarithmic function? example: Fine the derivative of g(x)= lnx^2

OpenStudy (anonymous):

find*

OpenStudy (turingtest):

It is about using the chain rule here.\[u(x)=x^2\]\[g(u)=\ln(u)\to g'(u)=\frac{1}{u}\]\[\frac{d}{dx}g(x)=g'(u)u'(x)\]So just plug back in terms of x for g'(u) and u'(x).